# Fiducial Inference for Random-Effects Calibration Models: Advancing Reliable Quantification in Environmental Analytical Chemistry

### v2

@@ 6 Data Analysis
We shall report the results from the analysis of both the cadmium and copper datasets mentioned earlier. One measurement from each laboratory was selected as the test sample for each concentration, while the remaining data were used to train the model. Since we are working with real-world datasets, it is essential to first evaluate the adequacy of the Rocke-Lorenzato model specified in (1) before performing the data analysis to estimate the unknown concentrations. To this end, the training data was used to fit the Rocke-Lorenzato model (1). Subsequently, 95% upper and lower calibration curves were constructed based on the fitted model, the majority of observed measurements fall within the specified limits, supporting the adequacy of the Rocke-Lorenzato model (1) for these datasets. A detailed goodness-of-fit assessments for the Rocke–Lorenzato calibration model on these datasets is presented in the supplementary material.
@@ 7 Discussion
The fiducial approach developed in this article is unique in that it allows estimates for a single laboratory or multiple laboratories analyzing split samples from the same environmental source to borrow strength from other laboratories for which calibration data using the same analytic methodology are available. The distinct advantage of this approach is that both the point estimate and confidence interval are more precise, and computationally it is easy to implement. The approach presented in this paper addresses a crucial challenge in analytical chemistry, namely the uncertainty in chemical measurements. Uncertainty is frequently underestimated in this field, leading to the erroneous belief that measurement estimates reflect true values. The techniques devised in this study facilitate the estimation of uncertainty related to the actual true concentration upon which these measurements are predicated. Given that measurements are typically collected from various instruments or analyses, and often across multiple laboratories, it is essential to consider these sources of variability when evaluating the overall uncertainty in the analytical process. Therefore, these variances should be factored into calculations that define the bounds for the underlying true concentration. In practical applications, when assessing a measurement against a regulatory threshold limit, or compliance standard (referred to as RT), we can compute a confidence interval for the true unknown concentration $X$. An exceedance is declared only if the lower confidence limit (LCL) exceeds the RT, providing 100$(1-\alpha)\%$ confidence that the RT has been surpassed. Thus, determination of the LCL with the prespecified confidence level is extremely important for comparisons to the RT, and our fiducial approach provides an LCL for this purpose.
@@
Supplementary Material discusses specific cases where traditional methods like MLE and MME become computationally infeasible due to limited sample sizes. It also includes comprehensive simulation results with complete tables and graphical analyses that demonstrate the performance of the fiducial approach covering both interval and point estimation, as well as detailed goodness‐of-fit assessments for the Rocke–Lorenzato calibration model on cadmium and copper datasets.

### v3

@@ 6 Data Analysis
We report the results from the analysis of both the cadmium and copper datasets mentioned earlier. One measurement from each laboratory was selected as the test sample for each concentration, while the remaining data were used to train the model. Since we are working with real-world datasets, it is essential to first evaluate the adequacy of the Rocke-Lorenzato model specified in (1) before performing the data analysis to estimate the unknown concentrations. To this end, the training data was used to fit the Rocke-Lorenzato model (1). Subsequently, 95% upper and lower calibration curves were constructed based on the fitted model, the majority of observed measurements fall within the specified limits, supporting the adequacy of the Rocke-Lorenzato model (1) for these datasets. A detailed goodness-of-fit assessments for the Rocke–Lorenzato calibration model on these datasets is presented in the supplementary material.
@@ 7 Discussion
[…]
Throughout this paper, we use the term “borrowing strength” in a purely descriptive sense to denote information sharing induced by the joint fiducial construction, and not to indicate an additional hierarchical (random-effects) model for the laboratory-specific calibration parameters. In the proposed fiducial method, borrowing strength occurs because the variance components $(\sigma_{\eta}^{2},\sigma_{\epsilon}^{2})$ are common across laboratories and are estimated by pooling information over all laboratories and concentrations via equations (6)-(7). This pooled uncertainty, together with lab-specific parameters $(\alpha_{i},\beta_{i})$, is then propagated into inference on unknown concentrations. As a result, concentration intervals are stabilized relative to analyses that treat each laboratory separately, while still allowing the calibration relationships to differ across laboratories through $(\alpha_{i},\beta_{i})$.
@@
Acknowledgements
The authors are grateful to two reviewers for their helpful comments and positive feedback. Their comments resulted in the clarification of some of the ideas and improved presentation of the results.
@@ Acknowledgements
Supplementary Material discusses specific cases where traditional methods like MLE and MME become computationally infeasible due to limited sample sizes. It also includes comprehensive simulation results with complete tables and graphical analyses that demonstrate the performance of the fiducial approach covering both interval and point estimation, as well as detailed goodness‐of-fit assessments for the Rocke–Lorenzato calibration model on cadmium and copper datasets. The final section of the supplementary material includes the cadmium and copper datasets used in the manuscript.